OpenStudy (anonymous):

Surface/area integrals. Evaluate the double integral zdS over the area of a cone, z=sqrt(x^2+y^2) between z=0 and z=1. So, the area I have to calculate is that of a circle with r=1. I transformed the variables to the spherical coordinates with x=cos theta and y = sin theta. Using the limits for r:[0,1] and theta:[0,2pi]... start of the area integral in next post

OpenStudy (anonymous):

\[\int\limits_{0}^{2 \pi}\int\limits_{0}^{1}r ^{2}\sqrt{1+?+?} dr d \theta\] Not sure what I need to put for the derivatives over r and theta

OpenStudy (anonymous):

So far I have \[\int\limits_{}^{}\int\limits_{s}^{}\sqrt{x^2+y^2}.\sqrt{1+4(x^2+y^2)} dxdy\] Then I plug in the transformation variables instead and their limits \[\int\limits_{0}^{2\pi}\int\limits_{0}^{1} r^2 \sqrt{1+4r^2} dr d \theta\] Now I have something I can try to solve.

ganeshie8 (ganeshie8):

z = f(x,y) = sqrt(x^2+y^2) f_x = x/sqrt(x^2+y^2) f_y = y/sqrt(x^2+y^2) (f_x)^2 + (f_y)^2 = 1

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

\[\large S = \iint\limits_D \sqrt{f_x^2 + f_y^2 +1} dxdt \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No, not just the root alone

ganeshie8 (ganeshie8):

\[\large S = \iint\limits_D \sqrt{1 +1} dxdy \]

ganeshie8 (ganeshie8):

\[\large S =\sqrt{2} \iint\limits_D 1 dxdy \]

OpenStudy (anonymous):

the root stands for the vector product. It's a normal.

OpenStudy (anonymous):

you still have to multiply it with the function itself

ganeshie8 (ganeshie8):

hmm work it in rectangular coordinates, its easy

OpenStudy (anonymous):

Hmm, I think I'm mixing it up with something else

ganeshie8 (ganeshie8):

This is the surface area of a surface `z = f(x,y)` formula : \[\large S = \iint\limits_D \sqrt{f_x^2 + f_y^2 +1} dxdy\]

OpenStudy (anonymous):

I have an example where z=1+2x+3y and the double integral is \[\int\limits_{}^{}\int\limits_{S}^{}x ^{2}yz dS\] Substitute the expression for z in the integral function and then you multiply with the squareroot of the partial derivatives you get out of the z expression. For the example in this post that would give \[\int\limits_{0}^{3}\int\limits_{0}^{2}x ^{2}y(1+2x+3y)\sqrt{2^{2}+3^{2}+1} dydx\]

OpenStudy (anonymous):

It's the function x squareroot of the partial derivatives. You use the squareroot not to replace z, but to get rid of the dz

OpenStudy (anonymous):

My source for the example I gave is from a youtube lecture by Dr Cris Tisdell

OpenStudy (anonymous):

My course \[\int\limits_{}^{}\int\limits_{\sum}^{} \left( F,d \sigma \right)=\int\limits_{}^{}\int\limits_{D}^{}\left( F(x(u,v)),x_u X x_v \right) dudv\]

OpenStudy (anonymous):

it's an inproduct

OpenStudy (anonymous):

I was wrong... that version again is for a vectorfield. You were correct. I just have to do the squareroot. I had this for the root alone \[\sqrt{1+\frac{ x² }{ x^2+y^2 }+\frac{ y^2}{ x^2+y^2 }}\] I replace it all with the spherical variables \[\sqrt{1+\frac{ r^2\cos^2\theta }{r^2}+\frac{ r^2\sin^2\theta }{r^2}} \] The r^2 eliminate each other, and then you have cos^2 + sin ^2 = 1 so the root becomes \[\sqrt{2}\] \[\int\limits_{0}^{2\pi}\int\limits_{0}^{1}r \sqrt{2} drd \theta = \sqrt{2}\]

OpenStudy (anonymous):

sorry the solution is \[\sqrt{2} \pi\]

ganeshie8 (ganeshie8):

nice :) we can do this using geometry also : \[\large \begin{align} S &= \iint\limits_D \sqrt{1+f_x^2 + f_y^2 } dxdy\\~\\ &=\iint\limits_D \sqrt{1+\frac{ x^2}{ x^2+y^2 }+\frac{ y^2}{ x^2+y^2 }} dxdy\\~\\ &=\iint\limits_D \sqrt{1+\frac{ x^2+y^2 }{ x^2+y^2 }} dxdy\\~\\ &=\iint\limits_D \sqrt{1+1} dxdy\\~\\ &= \sqrt{2} \iint\limits_D1dxdy\\~\\ &= \sqrt{2} \pi~~~\color{gray}{(\because \text{Area of unit circle = }\pi*1^2 = \pi)}\\~\\ \end{align}\]